OpenStudy (jollypikachu300):

HELP WILL FAN AND MEDAL find the value of each xpression write the answer in simplest form 4 1/2 - 1 1/25 + 6 2/5

OpenStudy (dmcdowell):

is this one big xpression or what

OpenStudy (jollypikachu300):

its a mixed fraction

OpenStudy (mathmale):

4 1/2 - 1 1/25 + 6 2/5. OK. A bit ambiguous, but readable. suggest that you write it like this:\[4\frac{ 1 }{ 2 }-1\frac{ 1 }{ 25 }+6\frac{ 2 }{ 5 }\]

OpenStudy (mathmale):

Next step will be to identify the lowest common denominator. What is it? Note that 5 divides into 25, but 2 does not.

OpenStudy (dmcdowell):

oh I see now.

OpenStudy (jollypikachu300):

I don't get it

OpenStudy (mathmale):

tHE GIVEN expression has 3 different denominators. List them, please.

OpenStudy (jollypikachu300):

idkhem